OpenStudy (yrivers36):

The revenue from renting 50 hotels rooms is $4945. If premium rooms cost $115 and regular rooms cost $80, find the number of each type of room rented. Im not sure how to set the equation.

OpenStudy (anonymous):

27 @ $115 and 23 @ $80

OpenStudy (anonymous):

let no of premium rooms be 'p' no of regular rooms be 'r' now total no of rooms = no of premium rooms + no of regular rooms = p +r from given data => p + r = 50 ...(1) so total revenue = revenue from renting premium hotels rooms + revenue from renting regular hotels rooms => 4945 = (no of premium rooms)(premium room cost) + (no of regular rooms)(regular room cost) => 4945 = 115p + 80r ...(2) we obtain the following equation by substituting the value of p from equation (1) in equation (2) 4945 = 115(50 - r) + 80r => 4945 = 5750 - 115r + 80r => 35r = 805 => r = 805/35 = 23 => p = 27

OpenStudy (anonymous):

let x equal the number of rooms at the $115 rate. Solve the following for x x 115+(50-x)80=4945

OpenStudy (yrivers36):

Thank both of you

OpenStudy (anonymous):

Your welcome.